### v2

# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

Abstract
We introduce softpick, a rectified, not sum-to-one, drop-in replacement for softmax in transformer attention mechanisms that eliminates attention sink and massive activations. Our experiments with 340M and 1.8B parameter models demonstrate that softpick achieves 0% sink rate consistently. The softpick transformers produce hidden states with significantly lower kurtosis and creates sparse attention maps. Quantized models using softpick outperform softmax on standard benchmarks, with a particularly pronounced advantage at lower bit precisions. Our analysis and discussion shows how softpick has the potential to open new possibilities for quantization, low-precision training, sparsity optimization, pruning, and interpretability. Our code is available at https://github.com/zaydzuhri/softpick-attention.
Figure 1: (left) Comparison between the attention maps when using softmax vs softpick and overall sink rate of the 340M models. (right) Largest hidden state activation per layer of the 340M models.
1 Introduction
The softmax function is widely used in statistics and particularly in machine learning as a way to normalize a vector of real numbers into a probability distribution. It has since been adopted as the de facto function to normalize the scores in the attention mechanism [3, 43] used in the transformer architecture [46]. The use of softmax in attention was a natural choice, since it represents the probability of a query "matching" to a key among many keys, which can be used to return values each weighted by that probability. However, our modern use of attention makes us question this intuition [29, 41]: why must they be non-zero probabilities that sum to one?
We continue to use softmax in practice because it is still the most capable normalization for attention. Softmax has great training stability with a neat Jacobian matrix resulting in a dense gradient, bounded Frobenius norm, and built-in regularization [37], and has been proven to be sufficiently expressive in its nonlinearity [7]. Despite the benefits, softmax brings several oddities detrimental to language modeling. Its sum-to-one nature forces out a now well-observed behavior named attention sink [51] where attention heads allocate a significant score towards a specific token, often the initial BOS token, which itself is semantically irrelevant [19]. These sinks are, as far as we know, harmless to downstream performance. However, another symptom of this same effect of softmax is the appearance of extreme hidden state activations, dubbed massive activations [44], which grow larger as the model scales [6]. These activations are especially problematic for quantization, with entire algorithms built around their existence. Additionally, most of the current low-precision training approaches try to work around these massive activations as they are a hassle to represent in low-bit precision. Removing these massive activations would open up a lot more possibilities.
We have learned to live with these quirks of attention, exactly because we have been hesitant to move away from softmax. In this paper, we propose the softpick function in an attempt to find a normalization function more fitting for attention and its use in transformers. The softpick function mitigates both attention sink and massive activations by reshaping the softmax function to avoid strict sum-to-one behavior and allowing rectified outputs. Our contributions are as follows:
  1. 1.
    We propose the softpick function as a drop-in replacement to softmax in attention.
  2. 2.
    We experiment by training 340M and 1.8B parameter transformer models from scratch and show how softpick compares to softmax in benchmarks and training behavior. We also show that quantized softpick models down to 2-bit precision outperform softmax.
  3. 3.
    We analyze the resulting attention maps, attention sinks, and massive activations, or lack thereof. We show that softpick returns more legible attention maps, a sink rate of 0%, and no massive activations in the hidden states.
  4. 4.
    We discuss the results and their implications for future work on quantization, low-precision training, sparsity, pruning, interpretability, and usage in other modalities. We also note an unsolved issue with softpick, in that it returns underscaled scores when used with long context, proposing possible solutions.
2 Background
The attention sink as a phenomenon was first given a name in [51] after observing that strong attention scores were given to tokens that were semantically irrelevant. These tokens can be any token that has minimal influence and appears often and consistently in training. Usually, the BOS token at the start of a sequence satisfies these conditions the best. Many works since then have observed, analyzed, and tried to explain when and why they appear. Barbero et al. [4] argue that attention sinks are a way for LLMs to avoid over-mixing, where the attention heads that have a sink act as approximate no-ops as to avoid representational collapse in the residual stream of deep models. Meanwhile, in a more empirical work, Gu et al. [19] investigate whether optimization, data distribution, loss function, or model architecture in language model pre-training is more correlated to the phenomenon. They conclude that attention sink happens for any transformer language model trained effectively on sufficient data. While optimizer hyperparameter choice can encourage a stronger sink, and loss function or data distribution can determine where the sink happens, ultimately the existence of the sink boils down to specifically the dependence of tokens’ attention scores to sum to one in the softmax normalization.
Massive activations are so called in [44] to describe extremely large values in the hidden states or residuals of the transformer. Very few token positions actually reach these large magnitudes in some feature dimensions of their hidden states, where normally they are relatively much smaller in magnitude. These massive activations are important to consider because they hinder quantization and low-precision training to some degree [13]. Importantly, massive activations have been determined to be caused directly by self-attention and the softmax function inside it that manifests itself in the form of attention sink [44]. Therefore, getting rid of attention sink also gets rid of massive activations.
Several of the studies mentioned above have attempted to mitigate attention sink to varying success. Gu et al. [19] used sigmoid attention without normalization to avoid attention sink. However, attention sinks still appear if normalization is reintroduced, leading them to the conclusion that softmax is indeed the issue. They also experimented with adding attention bias by appending an extra vector to the queries, keys, and/or values in the sequence dimension. This was also tested in [44]. However, this only works if the vectors are trainable, which means adding extra parameters to the model. They also seem to worsen the deeper the model, as the massive activations appear again in later layers. Most promising is the approach in [21] which tests an idea previously proposed in a blogpost [29], named softmax-1, adding a $+1$ in the denominator to relax the sum-to-one nature of softmax. However, this seems to only work well when used in conjuction with a custom optimizer, OrthoAdam, in order to avoid outlier activations. Owen et al. [33] further confirms that these KV bias techniques are not a reliable mitigation strategy for massive activations. We propose softpick as a drop-in solution to mitigate both attention sink and massive activations at once, without the need for extra parameters or custom optimizers.
3 Method
3.1 Softpick Function
Given a vector $\mathbf{x}\in\mathbb{R}^{N}$, we define the softpick function as
$Softpick(\mathbf{x})_{i}=\frac{ReLU(e^{x_{i}}-1)}{\sum_{j=1}^{N}|e^{x_{j}}-1|}$ (1)
with the rectified linear unit defined as $ReLU(x)=max(x,0)$ and the absolute value function written as $|x|$. Similar to softmax, we need a numerically safe version to use in practice. We define numerically safe softpick as
$Softpick(\mathbf{x})_{i}=\frac{ReLU(e^{x_{i}-m}-e^{-m})}{\sum_{j=1}^{N}|e^{x_{j}-m}-e^{-m}|+\epsilon}$ (2)
where $m$ is the maximum value inside $\mathbf{x}$ and $0<\epsilon\ll 1$ is to avoid division by zero in the case that all inputs are exactly zero. This function is a drop-in replacement to softmax in the attention mechanism:
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=Softpick\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (3)
3.2 Design Rationale
To understand the reasoning behind each component of the softpick function, it is best we go through a step-by-step recreation of the formula. We start from the vanilla softmax function.
${\color[rgb]{.5,.5,.5}Softmax(\mathbf{x})_{i}=\frac{e^{x_{i}}}{\sum_{j=1}^{N}e^{x_{j}}}}$ (4)
The reason why we still do not modify this function too much is that the characteristics of softmax in training are ideal for attention. We want to maintain as much of the Jacobian matrix as possible and keep the gradient norm bounded that way. The next step would be to induce "null attention", or allowing attention to output zeros. We can achieve this by moving down the exponential by one, making it output negative values $-1<y<0$ for negative inputs. We then rectify this output with ReLU.
${\color[rgb]{.5,.5,.5}Softmax(\mathbf{x})_{i}=\frac{{\color[rgb]{1,.5,0}ReLU(}e^{x_{i}}{\color[rgb]{1,.5,0}-1)}}{\sum_{j=1}^{N}{\color[rgb]{1,.5,0}ReLU(}e^{x_{j}}{\color[rgb]{1,.5,0}-1)}}}$ (5)
This function has several desirable properties. First, it allows zero-valued outputs, which has several benefits in attention such as sparsity optimizations and minimizing noise from the accumulation of unneeded values. Second, unlike softmax, the denominator does not grow unbounded as $N$ grows to infinity. This allows for a sharp output distribution regardless of sequence length, where only the positive inputs receive a score that sum to one. However, this is also its weakness. We experimented with this rectified-only softmax and found that it diverges from baseline softmax after a longer training period, and hypothesize that this is caused by heads "dying" and not being able to recover from outputting only zeros. This is obvious when we look at the Jacobian of the function, where negative inputs do not receive any gradients because they do not contribute to the output nor the denominator. Moreover, using this modification does not result in getting rid of attention sink. We go into more detail on this and show the experimental results of using a similar function in Appendix B. To fix the issue of negative inputs receiving minimal gradients, we allow them to contribute to the denominator.
${\color[rgb]{.5,.5,.5}Softmax(\mathbf{x})_{i}=\frac{ReLU(e^{x_{i}}-1)}{\sum_{j=1}^{N}{\color[rgb]{1,.5,0}Abs(}e^{x_{j}}-1{\color[rgb]{1,.5,0})}}}$ (6)
The absolute function $Abs(x)$ or $|x|$ allows us to rectify each element for summation, resulting in a positive sum, but does not alter the derivative of $e^{x}$ other than its sign, i.e. $\frac{d}{dx}|e^{x}|=-e^{x}$ for $x<0$. This is crucial to one of our goals, which is to maintain the desired properties from the Jacobian of softmax. This way, gradients can flow even when the input to the function is negative. More importantly, this asymmetry between the numerator and denominator removes the strict requirement to sum to one, which is the main cause of attention sink.
3.3 Derivative
Given $\mathbf{s}=Sofpick(\mathbf{x})\in\mathbb{R}^{N}$ as the output, the partial derivative or elements of the Jacobian matrix of the (numerically safe) softpick function can be written as such:
$\frac{\partial s_{i}}{\partial x_{j}}=\begin{cases}step(x_{i})\frac{e^{x_{i}-m}}{\Sigma}-sign(x_{j})\frac{e^{x_{j}-m}}{\Sigma}s_{i}&i=j\\ -sign(x_{j})\frac{e^{x_{j}-m}}{\Sigma}s_{i}&i\neq j\end{cases}$ (7)
with $step(x)$ being the step function that returns 00 if $x\leq 0$ and $1$ if $x>0$ and $sign(x)$ being the sign function that returns $-1$ if $x<0$ and $1$ if $x\geq 0$. Additionally, $\Sigma=\sum_{k=1}^{N}|e^{x_{k}-m}-e^{-m}|+\epsilon$ is the denominator of the softpick function. Although unlike softmax where only the outputs need to be kept for a naive backward pass implementation, the softpick derivative is nonetheless trivial for implementations that recompute the inputs, as is done in FlashAttention.
3.4 FlashAttention
The FlashAttention algorithm [10, 9] allows for online single-pass computation of attention, bypassing the quadratic memory requirement. Because the $ReLU(x)$ and absolute $|x|$ functions uphold the multiplicative property for positive-valued multipliers (such as $e^{x}$ for any $x$), deriving an online version of softpick is possible, thus making it compatible with FlashAttention. We provide the algorithms for the more recent FlashAttention-2 version of the forward pass and backward pass that use the softpick function instead of softmax in Appendix A.
4 Experiments
To validate the viability of the softpick function over softmax, we conduct experiments on Llama-style (pre-norm, RoPE, SwiGLU MLP) transformers trained from scratch. We train 4 models, two for each size: 340M and 1.8B parameters, each with a softmax variant and a softpick variant. The detailed training configuration for all models is shown in Table 1. We use the flash-linear-attention repository [53] and their Flame training framework (built on top of torchtitan [23]) because of their easily modifiable but fast triton [45] kernels. We train each 340M model on a total of 52 billion tokens and each 1.8B model on 104 billion tokens sampled from the 100B subset of the fineweb-edu [25] dataset. Each training run took approximately 18 hours for the 340M models and 116 hours for the 1.8B models on 8xH100 GPUs.
Table 1: Training configuration and hyperparameters for {340M, 1.8B} models.
Parameter Value Parameter Value
Hidden size {1024, 2048} Optimizer AdamW
Num. layers {24, 32} Learning rate {3e-4, 2e-4}
Num. heads {16, 32} LR schedule cosine
Num. KV heads {16, 32} Minimum LR (%) 10%
Sequence length 4096 Warmup steps {1000, 2000}
RoPE $\theta$ 10,000 Device batch size {16, 8}
Grad. accum. steps {1, 2}
Tied embeddings False Global batch size 128
Vocab size 32,000 Training steps {100000, 200000}
Softpick $\epsilon$ 1e-6 Max. gradient norm to clip 1.0
5 Results
Training Loss and Gradients
Figure 2: Training loss and gradient norm during training of 340M models.
We present the resulting training loss and gradient norm of the 340M model over 100,000 training steps (52B tokens) in Figure 2. We observed that the training loss of the softpick transformer tracks very closely to the vanilla softmax transformer, with only a 0.004 gap in the final training loss. This supports our design decision to maintain as much of the important components of the softmax function as possible. However, this does not hold at scale, as we observe a much larger gap of 0.12 in the training loss of the larger 1.8B model. This will be heavily reflected in the benchmark results. Tangentially, we observe that the total gradient norm of the models’ parameters during training exhibits noteworthy behavior. At the start up to a third of the way into training, the gradient norm of the softpick model is higher than softmax, but goes down and stays below softmax until the end, and does not rise at as fast of a rate as softmax. Also important to note is that the magnitude of some gradient norms of the softpick model are much higher than the gradient norm peaks of softmax. We use gradient clipping with a maximum norm of 1.0 and do not see any instability caused by these large gradients.
Benchmarks
Table 2: Comparison of softpick vs softmax 340M parameter model performance on downstream tasks. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 56.61 56.73 +0.12
Acc $\uparrow$ 60.35 61.11 +0.76
Lambada Acc $\uparrow$ 36.25 36.21 -0.04
Perplexity $\downarrow$ 30.33 28.63 -1.70
Piqa Acc Norm $\uparrow$ 66.59 66.49 -0.10
Acc $\uparrow$ 66.97 66.59 -0.38
Sciq Acc Norm $\uparrow$ 74.90 77.30 +2.40
Acc $\uparrow$ 83.20 83.60 +0.40
Wikitext Word Perplexity $\downarrow$ 23.85 24.32 +0.47
Table 3: Comparison of softpick vs softmax 1.8B parameter model performance on downstream tasks. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 67.21 62.04 -5.17
Acc $\uparrow$ 72.73 68.60 -4.13
Lambada Acc $\uparrow$ 49.43 43.51 -5.92
Perplexity $\downarrow$ 11.38 15.81 +4.43
Piqa Acc Norm $\uparrow$ 73.61 70.89 -2.72
Acc $\uparrow$ 73.78 71.27 -2.51
Sciq Acc Norm $\uparrow$ 86.40 80.40 -6.00
Acc $\uparrow$ 90.10 87.00 -3.10
Wikitext Word Perplexity $\downarrow$ 15.10 17.87 +2.77
We benchmark all models with general downstream tasks that are appropriate for the model sizes. We use the LM Evaluation Harness [18] to evaluate on these benchmarks: AI2 Reasoning Challenge (ARC-e) [8], Lambada [34], specifically the OpenAI variant, PIQA [5], and SciQ [49]. All 4 benchmarks report accuracy, with an addition of perplexity for Lambada. We also measure validation perplexity on Wikitext [27]. See Table 2 and Table 3 for full results. We observe equal if not slightly better performance from softpick compared to softmax on the 340M model. However, the 1.8B models show that softpick does not seem to scale well, at least when using the same hyperparameters as softmax. Overall accuracy on benchmarks and perplexity are worse at this scale.
Quantization
Figure 3: Quantization result of softmax vs softpick 340M parameter models.
Figure 4: Quantization result of softmax vs softpick 1.8B parameter models.
We also benchmark the effect of quantization on the trained models. We verified this using HQQ [2], BitsandBytes (BNB) [13, 14], and GPTQ [17]. These methods encompass different approaches, including those that utilize calibration data and those that do not. We use the same benchmarks as before. The results can be seen in Figure 3 and Figure 4 with more complete tables in Appendix C. At the 340M scale, when using softpick, the quantized models are consistently better than when softmax is used, and quantization is less damaging as the precision goes down. This becomes less clear at the 1.8B as the gap in performance of the base models is larger. However, the gap does close in the lower the precision. We discuss why this is in more detail in 7.1 after analyzing massive activations in 6.3.
6 Analysis
Table 4: Metrics for analysis of softmax and softpick in terms of sink rate, hidden activation distribution, and attention map sparsity. *Note that sparsity was calculating by counting exact zeros in the lower triangle, which means the softmax sparsity only comes from underflowed values.
Size Model Sink Rate % Hidden Activation Sparsity %
$\epsilon_{s}=0.2$ $\epsilon_{s}=0.3$ Kurtosis Min. Max.
340M Softmax 68.28 63.41 33510.81 -207.55 240.27 4.53*
Softpick 00.00 00.00 340.96 -45.03 36.21 99.34
1.8B Softmax 41.73 14.96 74456.81 -173.48 371.43 4.28*
Softpick 00.00 00.00 2193.27 -51.05 101.51 95.63
(a) 340M models
(b) 1.8B models
Figure 5: Box plots of the hidden state activations at some layers of the softmax (left, blue) and softpick (right, orange) models.
6.1 Attention Maps
Figure 6: Attention maps of softmax and softpick 340M models on 3 different input texts. Two heads are visualized: Head 4 of Layer 10 and Head 9 of Layer 21. See more attention maps in Appendix E.
To understand the difference that the softpick function makes compared to softmax, it is best to observe the attention maps directly. Attention maps are the matrices resulting from the computation of $\mathbf{A}=Soft\{max,pick\}(\mathbf{QK}^{T}/\sqrt{d_{k}})$. We prepare three text inputs as samples to run through the models and extract the attention maps. We provide some clear example attention maps from the 340M models in Figure 6 and show more attention maps of all models in Appendix E.
First, it is clear that softpick behaves very differently in regards to how it normalizes attention scores compared to softmax. Due to the rectified numerator in softpick, scores are functionally sparse and concentrated into specific spots and regions, surrounded by actual zero scores. Although more general heads that attend to a wider range of tokens do still exist in the softpick model. Second, and more important to our thesis, is that attention sinks are nowhere to be seen with softpick.
To see the sparsity effect of softpick, we calculate the sparsity ratio of the attention map for softpick and softmax by running 10 text samples through the model, collecting the attention maps, and averaging the the sparsity, which is calculated by counting the number of exact zeros. Note that since this is a causal language model, we only consider the lower triangular scores of the attention map. As shown in Table 4, our method results in attention scores with 99.34% sparsity on the 340M model and 95.63% on the 1.8B model. While softmax in theory cannot return exact zeros, we observe a non-zero sparsity because in practice floating point scores can underflow to zero.
6.2 Attention Sink
Attention sink can be identified by the large scores on the first column in the attention map, which manifests as a bright vertical line leftmost on the attention heatmaps. Notably, the heads that have heavy attention sinks i.e. large scores on the BOS token and close to none on other tokens with softmax, become either specifically tuned to only score specific trigger tokens, or are completely shut off when not needed. An example of this can be seen in Figure 6, on head 4 of the 10th layer. The softmax model has a clear attention sink and barely attends to other tokens, while the softpick model completely shuts off the head (all zero scores) on the first input, and attends to some specific tokens on the second and third input sample text. We believe this behavior is a more concrete version of the active-dormant heads in vanilla attention, a previously observed phenomenon [20]. We discuss this and its implications in 7.4.
In addition to visually observing attention sinks, we can use the sink rate metric proposed in [19], defined as $\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{1}^{l,h}>\epsilon_{s})$ which is the percentage of heads in the transformer that on average has an attention score above some threshold $\epsilon_{s}$ in the first token, where $\alpha_{1}^{l,h}$ is the average value of the first column in $\mathbf{A}$ at layer $l$ and head $h$. We calculate the sink rate of both models on 1000 samples from the validation set of SlimPajama [42] with $\epsilon_{s}=0.3$ (as suggested in [19]) and also $\epsilon_{s}=0.2$ and present the results in Table 4. The model using softpick has a sink rate of 0% for both thresholds and both model sizes, which means that no attention heads are overscoring the BOS token in any scenario, effectively eliminating attention sink completely. The effects of this can be seen when we analyze the hidden activations between transformer layers and the large values inside them.
6.3 Massive Activations
We analyze the effect of softpick on massive activations by examining the hidden state outputs after every transformer layer. We present some key metrics in Table 4 and box plots of the hidden state activation distribution of every some layers of the 340M and 1.8B models in Figure 5. These metrics are calculated by running 10 samples of text through the model and collecting all hidden state vectors to evaluate them all at once. The box plots visualize quite clearly how large the massive activations of the softmax model are, so much so as to making the boxes (Q1, median, and Q3) barely visible due to the large minimum and maximum values. This is especially prominent from the middle layers up to second to last layer. Meanwhile, the softpick models do not exhibit these massive activations. There is an exception in the last layer where it matches the softmax model, most likely due to the need to project back into the vocabulary space. This difference in activations is quite dramatic and can be seen clearly in the metrics as well. Kurtosis of all hidden state activations in the 340M model is significantly reduced from 33,510 to 340, a hundred-fold reduction. The minimum and maximum values are also reduced by an order of magnitude. Note that the minimum and maximum values of the softpick model are in fact from the last layer, all previous layers have even smaller activations.
7 Discussion and Implications
7.1 Quantization
Quantizing models presents challenges due to a scaling problem caused by large values within the activations [13]. Naive quantization of these values leads to inaccurate scaling since simply quantizing a vector containing such values leads to bias towards the maximum scores from the massive activation, which makes non-outliers underrepresented. A common solution is to use blockwise absmax quantization [13, 14]; however, this method requires storing numerous maximum values. QLoRA [14] mitigates this storage requirement through double quantization, which reduces the storage requirement for the maximum states for each block, but this introduces approximations, as the calculations are not exact, potentially reducing accuracy. It has been proposed that removing attention sinks, identified as significant outliers in activation, should help address this scaling problem. As shown in Figure 3, our 340M softpick model, outperforms softmax across various benchmarks and quantization techniques. The performance advantage of softpick becomes particularly pronounced when lower bit precision (lower qbit) is employed. Softpick also opens up the avenue for new quantization algorithms in the future that do not need to accommodate massive activations.
As shown in Table 4 and Figure 5, our method successfully eliminates these massive activations from the model. This makes activation quantization a promising area for further research as well. Previous work by Shen et al. [39] observed that these outliers hurt quantization performance, leading them to develop the TRIP method to reduce outlier impact. Similarly, Xiao et al. [52] discovered that the activation outliers are 100 times larger than the weight outliers, causing them to dominate the maximum measurements. With massive activations gone, quantization can maintain accuracy while using fewer bits, leading to better compression and faster inference. This makes our approach complementary to existing quantization techniques, potentially allowing for even more efficient model deployment.
7.2 Low-Precision Training
Low-precision training [48, 16, 35, 12] is an emerging topic, driven by recent hardware supporting FP8 [32] and FP4 [31] Tensor Cores, which offer higher FLOPS. Peng et al. [35] found that lower precision can be used for gradients and optimizers without significant accuracy loss or changes to hyperparameters. Fishman et al. [16] identified that sudden outliers in activations, potentially amplified by functions such as SwiGLU [38], disrupt the statistical assumptions needed for low-bit training and can cause divergence. They addressed this by introducing Smooth-SwiGLU which also enables them to quantize the optimizer’s second moment. To handle outliers and massive activations, Wang et al. [48] applied online Hadamard transformations to make activations more Gaussian-like, enabling aggressive low-bit training. Similarly, DeepSeek-AI et al. [12] created a fine-grained quantization scheme to reduce quantization errors caused by large activations.
Since outliers are a primary challenge for low-bit training [16], our method aims to produce activations without large activations or attention sink. We hypothesize that by removing this main obstacle, our approach will potentially allow simple low-bit training schemes to be used directly without issue, mitigating the divergence problem that occurs when massive activations emerge during training.
7.3 Sparsity
As discussed in section 6.1, softpick enables sparse attention. This sparsity offers potential for acceleration of inference through sparse GEMM kernels [36]. Furthermore, this sparsity allows computational savings by skipping operations involving zero-valued scores, conceptually similar to optimizations based on activation sparsity [24]. Specifically in our method, because the sparsity is guaranteed in the attention map (post-softpick), the matrix multiplication with the value matrix $\mathbf{V}$ only needs to be computed for the non-zero attention scores, potentially eliminating redundant calculations.
7.4 Active-Dormant Heads and Pruning
As outlined in Section 6.1, our observations indicate that some heads in the softpick models stay entirely inactive until activated by particular trigger tokens. As noted in [20], these heads are called active-dormant heads due to their behavior. In vanilla softmax transformers, these heads are identified by a large attention sink, a "drained" value state where the value projections are mostly zero, and a large residual activation. We believe that the active-dormant heads in softpick fill a similar purpose, but without the drawbacks that we see in softmax models. This opens up some avenues in head pruning [47, 40], as it is now much easier to identify dormant heads and prune them, without affecting the computation at all if that head would be outputting zero scores anyways.
7.5 Interpretability
Our attention map analysis shows that the attention maps produced by softpick are much more legible compared to those produced by softmax. Objectively, this is reflected in the sparsity metric discussed previously, but it can also be argued that the attention maps shown in Appendix E are visually more interpretable for softpick. This may have implications for interpretability techniques such as attention rollout [1].
7.6 Other Modalities
Attention sink and massive activation outliers have also been observed in visual transformers. Darcet et al. [11] identified artifacts within the feature maps of the Vision Transformers (ViT) [15] architecture. They addressed this problem by introducing register tokens designed to function as dedicated sinks. This is similar to adding attention bias in language models. Using softpick in ViTs might alleviate this issue and remove the need to use trainable register tokens that require extra parameters and computation. Attention sink has also been observed in video diffusion transformers [50] and streaming ASR models [22] where our function might help. Softpick could also potentially be useful in removing attention sinks in multimodal models.
7.7 Long Context and The Issue of Underscoring
Table 5: Passkey retrieval results of softmax vs softpick 340M models in terms of accuracy at certain sequence lengths.
Model Sequence Length
890 1983 3075 4167 4986
Softmax 95.5 97.0 73.0 70.5 0.0
Softpick 94.0 91.0 75.0 66.5 0.0
One of the potential benefits of sharper and sparse attention scores is the reduction of noise when accumulating values. Unlike softmax where every value must have a non-zero score, softpick can assign zero to unneeded values, theoretically leading to better retrieval for long context use. However, we have yet to see these benefits in our testing. We show some results on the passkey retrieval task [26] of both models on several sequence lengths in Table 5. Overall, softpick does not outperform softmax, at least not on this task.
We have a hypothesis as to why this is the case. The main issue is that softpick assigns smaller scores the more sparse the attention pattern and the longer the context is. Due to the nature of the denominator containing the absolute value function, if the long context has mostly negative scored tokens, those will significantly scale down the numerator on the few positive scored tokens. This can be seen in several of the attention maps in Appendix E, where the more retrieval-specific heads that only attend to singular tokens have a much smaller scale in scores (see the color bar of the heatmaps). This might result in too weak of a signal from the values, resulting in weak retrieval. We also hypothesize that this is one of the reasons why the larger 1.8B model with softpick does not perform as well as softmax in general.
One possible solution that we have considered is to scale up the input scores i.e. the query-key dot product result before going into softpick. This is an approach first proposed in Scalable-Softmax [30] and recently used in Llama 4 models [28]. However, we did not see significant improvement on passkey retrieval results and observed worse training loss and downstream task performance. We detail our experiments and results on this "scalable-softpick" approach in Appendix B. It remains an open problem to solve this issue and find a softmax alternative that scales better than softpick.
Acknowledgments
We would like to thank fal.ai and their grant program (https://fal.ai/grants) for providing us with the GPU hours needed to train all 340M and 1.8B models.
Appendix A FlashAttention Algorithms
Algorithm 1 FlashAttention-2 Forward Pass with Softpick
0:  Matrices $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d}$ in HBM, block sizes $B_{c}$, $B_{r}$, denominator epsilon $\epsilon$
1:   Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$ in to $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$ and $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$ each.
2:  Divide the output $\mathbf{O}\in\mathbb{R}^{N\times d}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, and divide the logsumexp $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3:  for $1\leq i\leq T_{r}$ do
4:      Load $\mathbf{Q}_{i}$ from HBM to on-chip SRAM.
5:      On chip, initialize $\mathbf{O}_{i}^{(0)}=(0)_{B_{r}\times d}\in\mathbb{R}^{B_{r}\times d},\ell_{i}^{(0)}=(0)_{B_{r}}\in\mathbb{R}^{B_{r}},m_{i}^{(0)}=(-\infty)_{B_{r}}\in\mathbb{R}^{B_{r}}$.
6:     for $1\leq j\leq T_{c}$ do
7:         Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to on-chip SRAM.
8:         On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
9:         On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$,$\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)})-\exp(-m_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{R}}_{i}^{(j)}=ReLU(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{A}}_{i}^{(j)}=|\tilde{\mathbf{P}}_{i}^{(j)}|\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{A}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
10:         On chip, compute $\mathbf{O}_{i}^{(j)}=\mathrm{diag}(e^{m_{i}^{(j-1)}-m_{i}^{(j)}})^{-1}\mathbf{O}_{i}^{(j-1)}+\tilde{\mathbf{R}}_{i}^{(j)}\mathbf{V}_{j}$.
11:     end for
12:     On chip, compute $\ell_{i}^{(T_{c})}=\ell_{i}^{(T_{c})}+\epsilon$.
13:     On chip, compute $\mathbf{O}_{i}=\mathrm{diag}(\ell_{i}^{(T_{c})})^{-1}\mathbf{O}_{i}^{(T_{c})}$.
14:     On chip, compute $L_{i}=m_{i}^{(T_{c})}+\log(\ell_{i}^{(T_{c})})$.
15:     Write $\mathbf{O}_{i}$ to HBM as the $i$-th block of $\mathbf{O}$.
16:     Write $L_{i}$ to HBM as the $i$-th block of $L$.
17:  end for
18:  Return the output $\mathbf{O}$ and the logsumexp $L$.
Algorithm 2 FlashAttention-2 Backward Pass with Softpick
0:  Matrices $\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{O},\mathbf{dO}\in\mathbb{R}^{N\times d}$ in HBM, vector $L\in\mathbb{R}^{N}$ in HBM, block sizes $B_{c}$, $B_{r}$.
1:  Divide $\mathbf{Q}$ into $T_{r}=\left\lceil\frac{N}{B_{r}}\right\rceil$ blocks $\mathbf{Q}_{1},\dots,\mathbf{Q}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $\mathbf{K},\mathbf{V}$ in to $T_{c}=\left\lceil\frac{N}{B_{c}}\right\rceil$ blocks $\mathbf{K}_{1},\dots,\mathbf{K}_{T_{c}}$ and $\mathbf{V}_{1},\dots,\mathbf{V}_{T_{c}}$, of size $B_{c}\times d$ each.
2:  Divide $\mathbf{O}$ into $T_{r}$ blocks $\mathbf{O}_{i},\dots,\mathbf{O}_{T_{r}}$ of size $B_{r}\times d$ each, divide $\mathbf{dO}$ into $T_{r}$ blocks $\mathbf{dO}_{i},\dots,\mathbf{dO}_{T_{r}}$ of size $B_{r}\times d$ each, and divide $L$ into $T_{r}$ blocks $L_{i},\dots,L_{T_{r}}$ of size $B_{r}$ each.
3:  Initialize $\mathbf{dQ}=(0)_{N\times d}$ in HBM and divide it into $T_{r}$ blocks $\mathbf{dQ}_{1},\dots,\mathbf{dQ}_{T_{r}}$ of size $B_{r}\times d$ each. Divide $\mathbf{dK},\mathbf{dV}\in\mathbb{R}^{N\times d}$ in to $T_{c}$ blocks $\mathbf{dK}_{1},\dots,\mathbf{dK}_{T_{c}}$ and $\mathbf{dV}_{1},\dots,\mathbf{dV}_{T_{c}}$, of size $B_{c}\times d$ each.
4:  Compute $D=\mathrm{rowsum}(\mathbf{dO}\circ\mathbf{O})\in\mathbb{R}^{d}$ (pointwise multiply), write $D$ to HBM and divide it into $T_{r}$ blocks $D_{1},\dots,D_{T_{r}}$ of size $B_{r}$ each.
5:  for $1\leq j\leq T_{c}$ do
6:     Load $\mathbf{K}_{j},\mathbf{V}_{j}$ from HBM to on-chip SRAM.
7:     Initialize $\mathbf{dK}_{j}=(0)_{B_{c}\times d},\mathbf{dV}_{j}=(0)_{B_{c}\times d}$ on SRAM.
8:     for $1\leq i\leq T_{r}$ do
9:        Load $\mathbf{Q}_{i},\mathbf{O}_{i},\mathbf{dO}_{i},\mathbf{dQ}_{i},L_{i},D_{i}$ from HBM to on-chip SRAM.
10:        On chip, compute $\mathbf{S}_{i}^{(j)}=\mathbf{Q}_{i}\mathbf{K}_{j}^{T}\in\mathbb{R}^{B_{r}\times B_{c}}$.
11:        On chip, compute $\mathbf{E}_{i}^{(j)}=\exp(\mathbf{S}_{ij}-L_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$, $\mathbf{P}_{i}^{(j)}=\mathbf{E}_{i}^{(j)}-\exp(-L_{i})\in\mathbb{R}^{B_{r}\times B_{c}}$, $\mathbf{R}_{i}^{(j)}=ReLU(\mathbf{P}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$.
12:        On chip, compute $\mathbf{dV}_{j}\leftarrow\mathbf{dV}_{j}+(\mathbf{R}_{i}^{(j)})^{\top}\mathbf{dO}_{i}\in\mathbb{R}^{B_{c}\times d}$.
13:        On chip, compute $\mathbf{dP}_{i}^{(j)}=\mathbf{dO}_{i}\mathbf{V}_{j}^{\top}\in\mathbb{R}^{B_{r}\times B_{c}}$, $\mathbf{dR}_{i}^{(j)}=step(\mathbf{S}_{i}^{(j)})\circ\mathbf{dP}_{i}^{(j)}\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise multiply or use where() function), $\mathbf{dA}_{i}^{(j)}=sign(\mathbf{S}_{i}^{(j)})\circ D_{i}\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise multiply or use where() function).
14:        On chip, compute $\mathbf{dS}_{i}^{(j)}=\mathbf{E}_{i}^{(j)}\circ(\mathbf{dR}_{i}^{(j)}-\mathbf{dA}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$.
15:        Load $\mathbf{dQ}_{i}$ from HBM to SRAM, then on chip, update $\mathbf{dQ}_{i}\leftarrow\mathbf{dQ}_{i}+\mathbf{dS}_{i}^{(j)}\mathbf{K}_{j}\in\mathbb{R}^{B_{r}\times d}$, and write back to HBM.
16:        On chip, compute $\mathbf{dK}_{j}\leftarrow\mathbf{dK}_{j}+{\mathbf{dS}_{i}^{(j)}}^{\top}\mathbf{Q}_{i}\in\mathbb{R}^{B_{c}\times d}$.
17:     end for
18:     Write $\mathbf{dK}_{j},\mathbf{dV}_{j}$ to HBM.
19:  end for
20:  Return $\mathbf{dQ},\mathbf{dK},\mathbf{dV}$.
Appendix B Other experiments
We present our negative results from other experiments related to softpick. First is using the following function inside attention, which we call "rectified-only softmax":
$\nu(x)$ =$=\begin{cases}x&x\geq 0\\ -\infty&x<0\end{cases}$
$Softmax(\mathbf{x})_{i}$ $=\frac{e^{\nu(x_{i})}}{\sum_{j=1}^{N}e^{\nu(x_{j})}}$
Second is the "scalable-softpick" approach, an attempt to improve long context performance taking reference from [30]:
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=Softpick\left(\frac{(s\,log(n))\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (8)
where $s$ is a per-head trainable parameter and $n$ is the sequence length. Note that in the training code, $n$ is a vector which assigns a different sequence length, the corresponding one, for each row i.e. token position. We train both models on the same setup and configuration as our two main 340M parameter models. We report the training loss and gradient norm plots in Figure 7, the benchmark results in Table 6, and passkey retrieval results in Table 7.
Figure 7: Training loss and gradient norm during training, including other experiments.
Table 6: Comparison of softmax, softpick, rectified-only softmax, and scalable-softpick performance on downstream tasks. $\uparrow$ = Higher is better, $\downarrow$ = Lower is better.
Task Metric Softmax Softpick Rectified-Only Scalable-Softpick
Arc Easy Acc Norm $\uparrow$ 56.31 56.61 51.52 55.89
Acc $\uparrow$ 60.61 60.82 57.66 60.98
Lambada Acc $\uparrow$ 36.35 36.21 29.05 33.17
Perplexity $\downarrow$ 30.33 28.67 53.77 34.49
Piqa Acc Norm $\uparrow$ 66.43 66.27 66.27 66.32
Acc $\uparrow$ 66.87 66.43 66.43 66.92
Sciq Acc Norm $\uparrow$ 75.10 77.20 68.80 75.00
Acc $\uparrow$ 83.30 83.60 78.80 83.70
Wikitext Word Perplexity $\downarrow$ 204.94 141.56 357.95 213.74
Table 7: Passkey retrieval results including the other experiments.
Model Sequence Length
890 1983 3075 4167 4986
Softmax 95.5 97.0 73.0 70.5 0.0
Softpick 94.0 91.0 75.0 66.5 0.0
Rectified-Only 95.5 68.0 57.5 24.5 0.0
Scalable-Softmax 98.0 96.9 87.0 55.0 0.0
Appendix C Full Result on Quantization
Table 8: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 4-bit 53.66 53.87 +0.21
Acc $\uparrow$ 4-bit 58.88 59.60 +0.72
Acc Norm $\uparrow$ 8-bit 55.51 56.36 +0.84
Acc $\uparrow$ 8-bit 59.81 60.82 +1.01
Lambada Acc $\uparrow$ 4-bit 33.63 31.73 -1.90
Perplexity $\downarrow$ 4-bit 37.59 39.25 +1.67
Acc $\uparrow$ 8-bit 35.53 35.86 +0.33
Perplexity $\downarrow$ 8-bit 31.22 28.74 -2.48
Piqa Acc Norm $\uparrow$ 4-bit 65.29 66.38 +1.09
Acc $\uparrow$ 4-bit 65.40 66.32 +0.92
Acc Norm $\uparrow$ 8-bit 66.32 66.32 +0.00
Acc $\uparrow$ 8-bit 67.36 66.38 -0.98
Sciq Acc Norm $\uparrow$ 4-bit 74.90 75.10 +0.20
Acc $\uparrow$ 4-bit 82.90 82.80 -0.10
Acc Norm $\uparrow$ 8-bit 75.20 77.40 +2.20
Acc $\uparrow$ 8-bit 83.40 83.50 +0.10
Wikitext Word Perplexity $\downarrow$ 4-bit 26.38 26.47 +0.09
8-bit 23.95 24.40 +0.46
Table 9: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 27.27 29.76 +2.48
Acc $\uparrow$ 2-bit 27.02 29.88 +2.86
Acc Norm $\uparrow$ 3-bit 48.57 51.81 +3.24
Acc $\uparrow$ 3-bit 53.83 56.65 +2.82
Acc Norm $\uparrow$ 4-bit 53.96 54.29 +0.34
Acc $\uparrow$ 4-bit 59.93 59.26 -0.67
Acc Norm $\uparrow$ 8-bit 56.31 56.36 +0.04
Acc $\uparrow$ 8-bit 60.31 60.94 +0.63
Lambada Acc $\uparrow$ 2-bit 0.04 0.43 +0.39
Perplexity $\downarrow$ 2-bit 930980.10 133928.18 -797051.92
Acc $\uparrow$ 3-bit 23.71 29.07 +5.36
Perplexity $\downarrow$ 3-bit 82.12 52.32 -29.80
Acc $\uparrow$ 4-bit 33.55 34.97 +1.42
Perplexity $\downarrow$ 4-bit 37.30 31.08 -6.21
Acc $\uparrow$ 8-bit 36.44 36.15 -0.29
Perplexity $\downarrow$ 8-bit 30.06 28.66 -1.41
Piqa Acc Norm $\uparrow$ 2-bit 48.42 53.59 +5.17
Acc $\uparrow$ 2-bit 52.12 54.52 +2.39
Acc Norm $\uparrow$ 3-bit 63.76 64.64 +0.87
Acc $\uparrow$ 3-bit 64.53 65.56 +1.03
Acc Norm $\uparrow$ 4-bit 66.10 66.43 +0.33
Acc $\uparrow$ 4-bit 65.89 66.59 +0.71
Acc Norm $\uparrow$ 8-bit 66.59 66.54 -0.05
Acc $\uparrow$ 8-bit 66.81 66.21 -0.60
Sciq Acc Norm $\uparrow$ 2-bit 29.80 35.80 +6.00
Acc $\uparrow$ 2-bit 27.80 34.80 +7.00
Acc Norm $\uparrow$ 3-bit 73.50 75.90 +2.40
Acc $\uparrow$ 3-bit 79.30 82.30 +3.00
Acc Norm $\uparrow$ 4-bit 73.40 76.40 +3.00
Acc $\uparrow$ 4-bit 81.00 82.70 +1.70
Acc Norm $\uparrow$ 8-bit 75.00 77.60 +2.60
Acc $\uparrow$ 8-bit 83.20 83.70 +0.50
Wikitext Word Perplexity $\downarrow$ 2-bit 4081.76 893.35 -3188.41
3-bit 34.57 31.53 -3.04
4-bit 25.69 25.53 -0.16
8-bit 23.84 24.31 +0.47
Table 10: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 26.68 36.36 +9.68
Acc $\uparrow$ 2-bit 26.81 37.54 +10.73
Acc Norm $\uparrow$ 3-bit 48.65 52.57 +3.91
Acc $\uparrow$ 3-bit 52.78 57.95 +5.18
Lambada Acc $\uparrow$ 2-bit 0.00 5.76 +5.76
Perplexity $\downarrow$ 2-bit 21857842.67 3741.81 -21854100.86
Acc $\uparrow$ 3-bit 26.18 32.14 +5.96
Perplexity $\downarrow$ 3-bit 76.81 41.85 -34.96
Piqa Acc Norm $\uparrow$ 2-bit 50.38 55.11 +4.73
Acc $\uparrow$ 2-bit 52.29 56.04 +3.75
Acc Norm $\uparrow$ 3-bit 62.40 65.78 +3.37
Acc $\uparrow$ 3-bit 63.55 66.00 +2.45
Sciq Acc Norm $\uparrow$ 2-bit 22.20 51.40 +29.20
Acc $\uparrow$ 2-bit 21.60 53.60 +32.00
Acc Norm $\uparrow$ 3-bit 71.60 72.10 +0.50
Acc $\uparrow$ 3-bit 80.50 80.70 +0.20
Wikitext Word Perplexity $\downarrow$ 2-bit 127870.20 2139.86 -125730.34
3-bit 35.52 30.24 -5.29
Appendix D Full Result on Quantization on 1.8B
Table 11: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 4-bit 66.75 58.84 -7.91
Acc $\uparrow$ 4-bit 71.93 67.05 -4.88
Acc Norm $\uparrow$ 8-bit 67.26 62.21 -5.05
Acc $\uparrow$ 8-bit 72.22 69.07 -3.16
Lambada Acc $\uparrow$ 4-bit 47.84 41.26 -6.58
Perplexity $\downarrow$ 4-bit 12.90 18.56 +5.66
Acc $\uparrow$ 8-bit 48.65 43.14 -5.51
Perplexity $\downarrow$ 8-bit 11.74 16.11 +4.36
Piqa Acc Norm $\uparrow$ 4-bit 72.14 70.57 -1.58
Acc $\uparrow$ 4-bit 72.63 71.00 -1.63
Acc Norm $\uparrow$ 8-bit 72.96 70.40 -2.56
Acc $\uparrow$ 8-bit 73.56 71.22 -2.34
Sciq Acc Norm $\uparrow$ 4-bit 86.90 79.40 -7.50
Acc $\uparrow$ 4-bit 90.90 86.90 -4.00
Acc Norm $\uparrow$ 8-bit 86.00 79.90 -6.10
Acc $\uparrow$ 8-bit 90.30 86.90 -3.40
Wikitext Word Perplexity $\downarrow$ 4-bit 16.17 18.97 +2.80
8-bit 15.20 17.94 +2.74
Table 12: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 28.62 34.60 +5.98
Acc $\uparrow$ 2-bit 29.67 33.88 +4.21
Acc Norm $\uparrow$ 3-bit 61.28 57.49 -3.79
Acc $\uparrow$ 3-bit 64.90 64.27 -0.63
Acc Norm $\uparrow$ 4-bit 65.74 61.87 -3.87
Acc $\uparrow$ 4-bit 70.20 68.73 -1.47
Acc Norm $\uparrow$ 8-bit 66.96 61.95 -5.01
Acc $\uparrow$ 8-bit 72.77 68.60 -4.17
Lambada Acc $\uparrow$ 2-bit 1.14 1.59 +0.45
Perplexity $\downarrow$ 2-bit 27256.52 18730.23 -8526.29
Acc $\uparrow$ 3-bit 40.71 36.77 -3.94
Perplexity $\downarrow$ 3-bit 20.58 25.33 +4.75
Acc $\uparrow$ 4-bit 49.64 40.95 -8.69
Perplexity $\downarrow$ 4-bit 11.43 17.94 +6.50
Acc $\uparrow$ 8-bit 49.60 43.41 -6.19
Perplexity $\downarrow$ 8-bit 11.35 15.82 +4.47
Piqa Acc Norm $\uparrow$ 2-bit 53.97 53.92 -0.05
Acc $\uparrow$ 2-bit 54.46 55.22 +0.76
Acc Norm $\uparrow$ 3-bit 71.00 70.51 -0.49
Acc $\uparrow$ 3-bit 69.91 70.18 +0.27
Acc Norm $\uparrow$ 4-bit 73.34 69.97 -3.37
Acc $\uparrow$ 4-bit 71.98 71.11 -0.87
Acc Norm $\uparrow$ 8-bit 73.83 70.67 -3.16
Acc $\uparrow$ 8-bit 73.72 71.27 -2.45
Sciq Acc Norm $\uparrow$ 2-bit 45.80 45.30 -0.50
Acc $\uparrow$ 2-bit 43.90 43.80 -0.10
Acc Norm $\uparrow$ 3-bit 82.20 78.30 -3.90
Acc $\uparrow$ 3-bit 88.00 85.50 -2.50
Acc Norm $\uparrow$ 4-bit 85.90 80.40 -5.50
Acc $\uparrow$ 4-bit 90.50 86.60 -3.90
Acc Norm $\uparrow$ 8-bit 86.40 80.40 -6.00
Acc $\uparrow$ 8-bit 90.00 87.20 -2.80
Wikitext Word Perplexity $\downarrow$ 2-bit 4660.08 336.22 -4323.86
3-bit 19.24 21.77 +2.52
4-bit 15.84 18.46 +2.62
8-bit 15.10 17.86 +2.77
Table 13: Comparison of softpick vs softmax performance for HQQ [2] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 67.21 61.99 -5.22
Acc $\uparrow$ 2-bit 72.64 68.60 -4.04
Acc Norm $\uparrow$ 3-bit 67.21 61.99 -5.22
Acc $\uparrow$ 3-bit 72.64 68.60 -4.04
Lambada Acc $\uparrow$ 2-bit 49.56 43.61 -5.96
Perplexity $\downarrow$ 2-bit 11.38 15.83 +4.45
Acc $\uparrow$ 3-bit 49.56 43.61 -5.96
Perplexity $\downarrow$ 3-bit 11.38 15.83 +4.45
Piqa Acc Norm $\uparrow$ 2-bit 73.45 70.89 -2.56
Acc $\uparrow$ 2-bit 73.78 71.38 -2.39
Acc Norm $\uparrow$ 3-bit 73.45 70.89 -2.56
Acc $\uparrow$ 3-bit 73.78 71.38 -2.39
Sciq Acc Norm $\uparrow$ 2-bit 86.30 80.40 -5.90
Acc $\uparrow$ 2-bit 90.20 87.10 -3.10
Acc Norm $\uparrow$ 3-bit 86.30 80.40 -5.90
Acc $\uparrow$ 3-bit 90.20 87.10 -3.10
Wikitext Word Perplexity $\downarrow$ 2-bit 15.10 17.86 +2.77
3-bit 15.10 17.86 +2.77
Appendix E More Attention Maps
To show the difference between attention maps produced by softmax and softpick, we run 3 sample text inputs through the models and extract the resulting attention maps. The sample text inputs are the following:
Input 1:
"The dominant sequence transduction models are based on complex recurrent or convolutional neural networks that include an encoder and a decoder. The best performing models also connect the encoder and decoder through an attention mechanism."
Input 2:
"According to all known laws of aviation, there is no way that a bee should be able to fly. Its wings are too small to get its fat little body off the ground. The bee, of course, flies anyway. Because bees don’t care what humans think is impossible."
Input 3:
"An idol whose dream is to become the owner of a fast food chain. Kiara is a phoenix, not a chicken or turkey (Very important). She burns brightly, working herself to the bone since she’ll just be reborn from her ashes anyway."
Figure 8: More attention maps of the softmax 340M model on sample text input 1.
Figure 9: More attention maps of the softpick 340M model on sample text input 1.
Figure 10: More attention maps of the softmax 340M model on sample text input 2.
Figure 11: More attention maps of the softpick 340M model on sample text input 2.
Figure 12: More attention maps of the softmax 340M model on sample text input 3.
Figure 13: More attention maps of the softpick 340M model on sample text input 3.
Figure 14: More attention maps of the softmax 1.8B model on sample text input 1.
Figure 15: More attention maps of the softpick 1.8B model on sample text input 1.
Figure 16: More attention maps of the softmax 1.8B model on sample text input 2.
Figure 17: More attention maps of the softpick 1.8B model on sample text input 2.
Figure 18: More attention maps of the softmax 1.8B model on sample text input 3.
Figure 19: More attention maps of the softpick 1.8B model on sample text input 3.